OpenStudy (calculusxy):

Mr. Daniels wants to cut a square in a wall for a new window. The window is 3 1/8 feet wide. The wall has a length a length of 15 3/8 feet. Mr. Daniels wants to position the window at the center of the wall. a) How far from each edge of the wall should the window be positioned?

OpenStudy (calculusxy):

@robtobey

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

@perl

OpenStudy (anonymous):

\[\frac{1}{2} \left(\left(15+\frac{3}{8}\right)-\left(3+\frac{1}{8}\right)\right)=\left(6+\frac{1}{8}\right) \text{ from each edge} \]

OpenStudy (calculusxy):

So it's 6 1/8 right?

OpenStudy (anonymous):

Yes, that is what I calculated. Becase the window in centered on the wall and the window has non zero width, then the widow's distance from each flanking wall is identical.

OpenStudy (calculusxy):

Yeah, that's what I had gotten too but I wanted to make sure.

OpenStudy (calculusxy):

Thank you :)

OpenStudy (anonymous):

Your welcome.